Hello, i need help solving this problem. I forgot how to do it so if someone can refresh my memory id appreciate it, thank you
Given f(x)=4-x and g(x)=X^2, find and simplify
A. f(-6)
b. g(x-6)
thank you once again.

Answer by Theo(13342)

f(x) = 4-x
g(x) = x^2

f(-6) = 4 - (-6) which is equal to 4 + 6 which is equal to 10.
you just replace the original argument with the new argument.
the original argument is x.
the new argument is -6.

g(x-6)= (x-6)^2 which is equal to x^2 - 12x + 36.
you just replace the original argument with the new argument.
the original argument is x.
the new argument is (x-6)